OpenStudy (justindrewbieber):

Divide 30 by half and add 10. What do you get ? #help

OpenStudy (birdy258):

12

OpenStudy (faiqraees):

70

OpenStudy (alexvila95):

25

OpenStudy (faiqraees):

dividing by half actually means multiplying by 2

OpenStudy (birdy258):

30 -half wich is 15 so yea 25

OpenStudy (faiqraees):

divide 30 by half 30 / 0.5 30 / 1 /2 60 add 10 70

OpenStudy (samsungfanboy):

Justin Bieber has entered the site. This is war

OpenStudy (faiqraees):

google it right now 30 divide by 0.5 is 60

OpenStudy (luigi0210):

@birdy258 could be right too ;) "Divide \(\color{red}{30} \) by \(\color{red}{half} \)" Could mean 30/15

imqwerty (imqwerty):

its known that 30/2 is 15 and you say that 30/0.5 is also 15 so doesn't this prove that 0.5=2 :) which is false

OpenStudy (faiqraees):

just see for yourself what calculation it doing it is diving 30/1 by 2

OpenStudy (faiqraees):

@Luigi0210 it doesnt say its half it just say half

OpenStudy (luigi0210):

But I wouldn't know ')

OpenStudy (faiqraees):

divide 30 by half 30/0.5 60 (did with a calculator) added 10 70

imqwerty (imqwerty):

\(\Large \frac{30}{0.5}\) multiply numerator and denominator with 2 \(\Large \frac{30 \times 2}{0.5 \times 2}\) =\(\Large \frac{60}{1}\)